what is y2 - y1 / x2 - x1 I need to know for the math test in 3 hours. PLEASE HELP ME.
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Perhaps, if you have a test on this 3 hours from the time you posted the message, and you don't yet know what this is, then you should consider dropping the class.

And when you take the class again, pay attention and keep up with the class so you don't have to ask for help like this.


Answer by ikleyn(52781)

.

It may have different names, depending on the context, which we do not know from your post.

For example, in one context, it is a slope of a straight line in a coordinate plane.

In other context, it is the first difference of a function y = y(x).

In the third context, it can be a formula to estimate the first derivative of the function y = y(x).

In the fourth context, it can be the average rate of changing function y = y(x).


At this forum, you can obtain comprehensive, precise, brilliant first class solutions and explanations to many problems;
but do not postpone your question/questions to the last 3 hours before the test: do it in advance.